OpenStudy (anonymous):

factor completely: 3x^2+14x+8

OpenStudy (anonymous):

3x^2+12x-2x+8 --> [3x(x+4)-2(x+4) ---> (3x-2)(x+4)

OpenStudy (anonymous):

I factored by grouping.

OpenStudy (anonymous):

wait no that might be wrong. o_o

OpenStudy (anonymous):

oh it's (3x+2)(x+4)

OpenStudy (anonymous):

3x² - 14x + 8 Write 2 sets of parentheses. (_____)(_____) = Look at the first term's factors. 3x² = 3x * x Write 3x and x at the beginning of each set of parentheses. (3x_____)(x_____) = Look at the last term's factors. You want the pair that will add up to the -14x when multiplied by the 3x and x you already have. 1 * 8 ===> (3x * 1) + (1x * 8) = 3x + 8x = 11x <== wrong pair 2 * 4 ===> (3x * 2) + (1x * 4) = 6x + 4x = 10x <== wrong pair 4 * 2 ===> (3x * 4) + (1x * 2) = 12x + 2x = 14x <== wrong pair 8 * 1 ===> (3x * 8) + (1x * 1) = 24x + 1x = 25x <== wrong pair -1 * -8 ===> (3x * -1) + (1x * -8) = -3x + -8x = -11x <== wrong pair -2 * -4 ===> (3x * -2) + (1x * -4) = -6x + -4x = -10x <== wrong pair -4 * -2 ===> (3x * -4) + (1x * -2) = -12x + -2x = -14x <== right pair -8 * -1 ===> (3x * -8) + (1x * -1) = -24x + -1x = -25x <== wrong pair -4 * -2 is the right pair. Since -4 is paired with 3x and 3x is on the outside, write -4 on the outside. (3x_____)(x - 4) = Since -2 is paired with x and x is on the inside, write -2 on the inside. (3x - 2)(x - 4) ANSWER: (3x - 2)(x - 4) CHECK: (3x - 2)(x - 4) = 3x(x) + 3x(-4) - 2(x) - 2(-4) = 3x² - 12x - 2x + 8 = 3x² - 14x + 8

OpenStudy (anonymous):

MINE IS RIGHT I FOILED! >_<

OpenStudy (anonymous):

if you FOIL mine, you'll see -.- his ends up being 3x^2-14x+8

OpenStudy (anonymous):

(3x+2)(x+4) foiled is 3x^2+14x+8

OpenStudy (anonymous):

thank you